# Simulating the Real World: A Unified Survey of Multimodal Generative Models

### v1

@@ Abstract
Understanding and replicating the real world is a critical challenge in Artificial General Intelligence (AGI) research. To achieve this, many existing approaches, such as world models, aim to capture the fundamental principles governing the physical world, enabling more accurate simulations and meaningful interactions. However, current methods often treat different modalities, including 2D (images), videos, 3D, and 4D representations, as independent domains, overlooking their interdependencies. Additionally, these methods typically focus on isolated dimensions of reality without systematically integrating their connections. In this survey, we present a unified survey for multimodal generative models that investigate the progression of data dimensionality in real-world simulation. Specifically, this survey starts from 2D generation (appearance), then moves to video (appearance+dynamics) and 3D generation (appearance+ geometry), and finally culminates in 4D generation that integrate all dimensions. To the best of our knowledge, this is the first attempt to systematically unify the study of 2D, video, 3D and 4D generation within a single framework. To guide future research, we provide a comprehensive review of datasets, evaluation metrics and future directions, and fostering insights for newcomers. This survey serves as a bridge to advance the study of multimodal generative models and real-world simulation within a unified framework.
@@ 1 Introduction
For decades, the research community has aspired to develop systems that encapsulate the fundamental principles of the physical world, a cornerstone in the journey towards Artificial General Intelligence (AGI) [1]. Central to this endeavor is simulating the real world with machines, aiming to capture the complexities of reality through the lens of multimodal generative models. The resultant world simulator holds the promise of advancing the understanding of real world, unlocking transformative applications such as virtual reality [2], games [3], robotics [4] and autonomous driving [5].
@@
The term “world simulator” was first introduced by Ha David [6], drawing an analogy to the concept of a mental model [7] in cognitive science. Building on this perspective, modern researches [8] formulate the simulator as an abstract framework that enables intelligent systems to simulate the real world through multimodal generative models. These models encode visual contents and spatial-temporal dynamics of the real world into compact representations. As geometry, appearance, and dynamics jointly contribute to the realness of generated contents, these three aspects are widely investigated by the community [9]. Traditional real-world simulation methods have long relied on graphics techniques that incorporate geometry, texture, and dynamics. Specifically, geometry and texture modeling [10] are employed to create the objects while methods like keyframe animation [11] and physics-based simulation [12] are adopted to simulate the movement and behavior of objects over time.
@@
Despite great progress, these traditional methods often require extensive manual designs, heuristic rule definitions, and computationally expensive processing, limiting their scalability and adaptability to diverse scenarios. Recently, learning-based approaches, particularly multimodal generative models, have revolutionized content creation by providing a data-driven approach to realistic simulations. These approaches reduce the reliance on manual efforts, improve generalization across tasks, and enable intuitive interactions between human and models. For example, Sora [13] has garnered significant attention for its realistic simulation capabilities, demonstrating an early-stage understanding of physical laws. The emergence of such generative models introduces new perspectives and methodologies, addressing the limitations of traditional methods by reducing the need for extensive manual design and computationally expensive modeling while enhancing adaptability and scalability in diverse simulation scenarios.
@@
Though existing generative models offer powerful techniques for synthesizing realistic content in distinct data dimensions, the real world exhibits inherently high-dimensional complexity, and a comprehensive review that systematically integrates these advancements across different dimensions is still absent. This survey aims to bridge this gap by unifying the study of real-world simulation through from the perspective of data dimension growth, as illustrated in Figure 1. Specifically, we start from 2D generation (appearance only) and then extend it to video and 3D generation by incorporating dynamics and geometry dimensions, respectively. Finally, we culminates in 4D generation by integrating all dimensions.
@@
In summary, this survey makes three key contributions. First, it provides a systematic review of methods for real world simulation from the perspective of data dimension growth through the lens of multimodal generative models. To the best of our knowledge, this is the first survey that unifies the study of 2D, video, 3D, and 4D generation, offering a structured and comprehensive overview of this research area. Second, it surveys the commonly used datasets, their properties, and the corresponding evaluation metrics from various perspectives. Third, it identifies open research challenges, aiming to provide guidance for further exploration in this field.
@@
We envision this survey as a resource that provides valuable insights for newcomers and fosters critical analysis among experienced researchers. The survey is organized as follows. Section 2 introduces the foundational concepts of deep generative models. Section 3 presents four key paradigms: 2D, video, 3D , and 4D generation. Section 4 reviews datasets and evaluation metrics for these paradigms. Finally, Section 5 outlines future directions, and Section 6 concludes the survey.
2 Preliminaries
@@
Deep generative models learn complicated and high-dimensional data distribution with the aid of deep neural networks. Denote the data sample as $\mathbf{x}$ and its distribution as $p_{data}(\mathbf{x})$, the objective of deep generative models is to approximate the $p_{data}(\mathbf{x})$ with $p_{\theta}(\mathbf{x})$, where $\theta$ is the parameter of models. In this section, we briefly review several mainstream generative models (Table I), including generative adversarial networks (GANs) [14], variational autoencoders (VAEs) [15], autoregressive models (AR Models) [16], normalizing flows (NFs) [17], and diffusion models [18].
2.1 Generative Adversarial Networks (GANs)
@@
$\mathbb{E}_{\mathbf{z}\sim p_{\mathbf{z}}(\mathbf{z})}[\log(1-D(G(\mathbf{z};\theta_{g});\theta_{d}))]$
@@
GANs [14] Flexible Unstable training
VAEs [15] Data compression 1) Posterior collapse 2) Blur
AR Models [16] Explicit density 1) Markov assumption
@@
NFs [17] Explicit density Limited capacity
Diffusion Models [18] High-quality samples Expensive computation
2.2 Variational Autoencoders (VAEs)
@@
[…]
where $p(\mathbf{z})$ is a prior distribution of $\mathbf{z}$, and $p_{\theta}(\mathbf{x}|\mathbf{z})$ is the distribution of $\mathbf{x}$ conditioned on $\mathbf{z}$. However, since this integration is usually intractable, VAEs maximize the lower bound of $\log p_{\theta}(\mathbf{x})\geq-\mathtt{KL}(q_{\theta}(\mathbf{z}|\mathbf{x})||p_{\theta}(\mathbf{z}))+\mathbb{E}_{q_{\theta}(\mathbf{z}|\mathbf{x})}[\log p_{\theta}(\mathbf{x}|\mathbf{z})]$, where $\mathtt{KL}(q_{\theta}(\mathbf{z}|\mathbf{x})||p_{\theta}(\mathbf{z}))$ is the KL divergence between $q_{\theta}(\mathbf{z}|\mathbf{x})$ and $p_{\theta}(\mathbf{z})$, and $\mathbb{E}_{q_{\theta}(\mathbf{z}|\mathbf{x})}[\log p_{\theta}(\mathbf{x}|\mathbf{z})]$ is computed by the Stochastic Gradient Variational Bayes estimator [15].
2.3 Autoregressive Models (AR Models)
Autoregressive Models (AR) models factorize $p_{\theta}(\mathbf{x})$ as the product of conditional probabilities to alleviate the difficulty of modeling multivariate joint probability,
$p_{\theta}(\mathbf{x})=p(x_{1},...,x_{d})=\prod_{i=1}^{d}p_{\theta}(x_{i}|x_{1},...,x_{i-1}),$ (3)
where $d$ is the sequence length.
Recently, numerous efforts have been made to employ AR models to model pixels in an image sequentially [19, 20, 21].
2.4 Normalizing Flows (NFs)
@@
where $g^{-1}(\cdot)$ is the inverse of $g(\cdot)$, and $\mathtt{J}(g(\mathbf{z}))=\frac{\partial g}{\partial\mathbf{z}}$ is the Jacobian of $\mathbf{g}$. NFs construct arbitrarily complicated non-linear invertible functions by compositing a set of $N$ bijective functions and define $g=g_{N}\circ\dots\circ g_{1}$.
2.5 Diffusion Models
@@
[…]
Since $Z_{\theta}$ is intractable to compute, the diffusion models learn the score function $\nabla_{\mathbf{x}}\log p_{\theta}(\mathbf{x})=-\nabla_{\mathbf{x}}s_{\theta}(\mathbf{x})$ instead.
@@
3 Paradigms
This section presents methods for simulating the real world from the perspective of data dimension growth. It begins with 2D generation (Section 3.1) for appearance modeling and then move to video generation (Section 3.2) and 3D generation (Section 3.3) by incorporating dynamics and geometry dimensions. Finally, by integrating all these three dimensions, recent advances in 4D generation (Section 3.4) are presented.
3.1 2D Generation
@@ 3.1.1 Algorithms
DeepFloyd IF [30] is celebrated for its exceptional photorealism and advanced language understanding. This system is modular, featuring a static text encoder and three sequential pixel diffusion modules. Initially, the base model creates 64$\times$64 pixel images from textual descriptions. These images are then enhanced to 256$\times$256 pixels and further to 1024$\times$1024 pixels by two super-resolution models. Each phase utilizes a static text encoder derived from the T5 [25] transformer to generate text embeddings, which are subsequently processed by a U-Net architecture with integrated cross-attention and attention pooling mechanisms.
Stable Diffusion (SD) [31], also known as Latent Diffusion Model (LDM), enhances training and inference efficiency on limited computational resources while producing high-quality and diverse images. The denoising process takes place in the latent space of pre-trained autoencoders, which map images into a spatial latent space. The underlying U-Net architecture is augmented with a cross-attention mechanism to model the conditional distribution, which can include text prompts, segmentation masks, and more. It used BERT tokenizer [24] as text encoding and trained on the LAION-400M [32] dataset to generate images at a resolution of 256$\times$256 (with a latent resolution of 32$\times$32). Building on Stable Diffusion, SDXL [33] employs a U-Net backbone that is three times larger. It introduces additional attention blocks and a larger cross-attention context by utilizing a second text encoder. Additionally, SDXL includes a refinement model that enhances the visual fidelity of samples generated by SDXL through a post-hoc image-to-image technique.
@@
3.2 Video Generation
Text-to-video generation models adapt text-to-image frameworks to handle the additional dimension of dynamics in the real world. We classify these models into three categories based on different generative machine learning architectures. Figure 2 summarizes recent text-to-video generation techniques. For further surveys, readers can refer to more detailed reviews [62, 63] in this subfield.
3.2.1 Algorithms
@@
3.2.2 Applications
@@
(3) Human animation in videos has gained significance in video generation, which plays a pivotal role in the world simulator, as discussed in section 3.2.1. This is particularly important because humans are the most essential participants in the real world, making their realistic simulation crucial. Thanks to the early success in generative models, there are some representative works [37, 89, 90] that introduce generative adversarial networks (GAN) [14] for animating humans in videos. Despite these progresses, the most essential problem in human video animation is the visual fidelity of the generated videos. ControlNet [91] and HumanSD [92] are plug-and-play methods for animating humans referring to poses based on a foundation text-to-image model, like Stable Diffusion [31]. Additionally, to resolve the generalization issue of these methods, animate-anyone [93] proposes a ReferenceNet to maintain more spatial details of the reference video and pushes generation quality in the wild to a new milestone. There are some follow-up works [94, 95] of animate-anyone trying to simplify the training architecture and costs. Besides, as the in-depth study of geometry and textures in the computer graphics area, there are some works that introduce 3D modeling into human video animation. Liquid Warping GAN [96], CustomHuman [97] and LatentMan [98] are early attempts introducing 3D human prior into the generation loop. The latest progress, MIMO [99], explicitly models the charter, 3D motion, and scene respectively to drive human animation in the wild. These methods with or without 3D priors make a great step to introduce humans into the loop of the world simulator.
@@
3.3 3D Generation
3D generation focuses on both geometry and appearance to better simulate real-world scenarios. In this section, we explore various 3D representations and generation algorithms, providing a structured overview of recent advancements. Specifically, we categorize 3D generation methods based on their input modalities, including Text-to-3D Generation, which directly synthesizes 3D content from textual descriptions, Image-to-3D Generation, which introduces image constraints to refine text-driven outputs, and Video-to-3D Generation, which leverages video priors for more consistent 3D generation. A chronological summary of these advancements is presented in Figure 4, while Table II offers a comprehensive comparison of cutting-edge methods. Notably, several approaches span multiple categories, demonstrating the versatility of modern 3D generation techniques.
3.3.1 3D Representation
@@
3.3.2 Algorithms

### v2

@@ Abstract
Understanding and replicating the real world is a critical challenge in Artificial General Intelligence (AGI) research. To achieve this, many existing approaches, such as world models, aim to capture the fundamental principles governing the physical world, enabling more accurate simulations and meaningful interactions. However, current methods often treat different modalities, including 2D (images), videos, 3D, and 4D representations, as independent domains, overlooking their interdependencies. Additionally, these methods typically focus on isolated dimensions of reality without systematically integrating their connections. In this survey, we present a unified survey for multimodal generative models that investigate the progression of data dimensionality in real-world simulation. Specifically, this survey starts from 2D generation (appearance), then moves to video (appearance+dynamics) and 3D generation (appearance+ geometry), and finally culminates in 4D generation that integrate all dimensions. To the best of our knowledge, this is the first attempt to systematically unify the study of 2D, video, 3D, and 4D generation within a single framework. To guide future research, we provide a comprehensive review of datasets, evaluation metrics, and future directions to foster insights for newcomers. This survey serves as a bridge to advance the study of multimodal generative models and real-world simulation within a unified framework.
@@ I Introduction
FOR decades, the research community has aspired to develop systems that encapsulate the fundamental principles of the physical world, a cornerstone in the journey towards Artificial General Intelligence (AGI) [1]. Central to this endeavor is simulating the real world with machines, aiming to capture the complexities of reality through the lens of multimodal generative models. The resultant world simulator holds the promise of advancing the understanding of the real world, unlocking transformative applications such as virtual reality [2], games [3], robotics [4], and autonomous driving [5].
@@
The term “world simulator” was first introduced by Ha David [6], drawing an analogy to the concept of a mental model [7] in cognitive science. Building on this perspective, modern researchers [8] formulate the simulator as an abstract framework that enables intelligent systems to simulate the real world through multimodal generative models. These models encode visual contents and spatial-temporal dynamics of the real world into compact representations. As geometry, appearance, and dynamics jointly contribute to the realness of generated contents, these three aspects are widely investigated by the community [9]. Traditional real-world simulation methods have long relied on graphics techniques that incorporate geometry, texture, and dynamics. Specifically, geometry and texture modeling [10] are employed to create the objects, while methods like keyframe animation [11] and physics-based simulation [12] are adopted to simulate the movement and behavior of objects over time.
@@
Despite great progress, these traditional methods often require extensive manual designs, heuristic rule definitions, and computationally expensive processing, limiting their scalability and adaptability to diverse scenarios. Recently, learning-based approaches, particularly multimodal generative models, have revolutionized content creation by providing a data-driven approach to realistic simulations. These approaches reduce the reliance on manual efforts, improve generalization across tasks, and enable intuitive interactions between humans and models. For example, Sora [13] has garnered significant attention for its realistic simulation capabilities, demonstrating an early-stage understanding of physical laws. The emergence of such generative models introduces new perspectives and methodologies, addressing the limitations of traditional methods by reducing the need for extensive manual design and computationally expensive modeling while enhancing adaptability and scalability in diverse simulation scenarios.
@@
Though existing generative models offer powerful techniques for synthesizing realistic content in distinct data dimensions, the real world exhibits inherently high-dimensional complexity, and a comprehensive review that systematically integrates these advancements across different dimensions is still absent. This survey aims to bridge this gap by unifying the study of real-world simulation from the perspective of data dimension growth, as illustrated in Figure 1. Specifically, we start from 2D generation (appearance only) and then extend it to video and 3D generation by incorporating dynamics and geometry dimensions, respectively. Finally, we culminate in 4D generation by integrating all dimensions.
@@
In summary, this survey makes three key contributions. First, it provides a systematic review of methods for real-world simulation from the perspective of data dimension growth through the lens of multimodal generative models. To the best of our knowledge, this is the first survey that unifies the study of 2D, video, 3D, and 4D generation, offering a structured and comprehensive overview of this research area. Second, it surveys the commonly used datasets, their properties, and the corresponding evaluation metrics from various perspectives. Third, it identifies open research challenges, aiming to guide further exploration in this field.
Previous surveys on generation models mainly concentrates on text-to-image, text-to-video, and text-to-3D generation separately without comprehensively studying their relationships. In contrast, this survey seeks to provide a more integrated perspective on multimodal generative models by tracing how generative models have evolved from handling appearance alone (2D generation), to incorporating dynamics (video generation) and geometry (3D generation), and ultimately to integrating appearance, dynamics, and geometry in 4D generation. This dimensional perspective aims to bridge previously isolated research areas and highlight common challenges and opportunities across them.
@@
We envision this survey as a resource that provides valuable insights for newcomers and fosters critical analysis among experienced researchers. The survey is organized as follows. Section II introduces the foundational concepts of deep generative models. Section III presents four key paradigms: 2D, video, 3D , and 4D generation. Section IV reviews datasets and evaluation metrics for these paradigms. Finally, Section V outlines future directions, and Section VI concludes the survey.
II Preliminaries
@@
Deep generative models learn complicated and high-dimensional data distributions with the aid of deep neural networks. Denote the data sample as $\mathbf{x}$ and its distribution as $p_{data}(\mathbf{x})$, the objective of deep generative models is to approximate the $p_{data}(\mathbf{x})$ with $p_{\theta}(\mathbf{x})$, where $\theta$ is the parameter of models. In this section, we briefly review several mainstream generative models (Table I), including generative adversarial networks (GANs) [14], variational autoencoders (VAEs) [15], autoregressive models (AR Models) [16], normalizing flows (NFs) [17], and diffusion models [18].
II-A Generative Adversarial Networks (GANs)
@@
$\mathbb{E}_{\mathbf{z}\sim p_{\mathbf{z}}(\mathbf{z})}[\log(1-D(G(\mathbf{z};\theta_{g});\theta_{d}))]$ .
There are several challenges associated with GAN training. For example, the Nash equilibrium may not always exist [19] or be hard to achieve [20], resulting in unstable training. Another problem is mode collapse, where the generator produces only specific types of samples with low diversity [20, 21].
@@
GANs [14] 1) Flexible 1) Unstable training
2) Efficient inference 2) Lack diversity
3) Mode collapse
VAEs [15] 1) Data compression 1) Posterior collapse
2) Blur
AR Models [16] 1) Explicit density 1) Markov assumption
@@
NFs [17] 1) Explicit density 1) Limited capacity
2) Restricted architecture
Diffusion Models [18] 1) High-quality samples 1) Expensive computation
2) Learn complex distribution
II-B Variational Autoencoders (VAEs)
@@
$p_{\theta}(\mathbf{x})=\int p_{\theta}(\mathbf{x}|\mathbf{z})p_{\theta}(\mathbf{z})d\mathbf{z},$ (2)
where $p_{\theta}(\mathbf{z})$ is a prior distribution of $\mathbf{z}$, and $p_{\theta}(\mathbf{x}|\mathbf{z})$ is the distribution of $\mathbf{x}$ conditioned on $\mathbf{z}$. However, since this integration is usually intractable, VAEs maximize the lower bound of $\log p_{\theta}(\mathbf{x})\geq-\mathtt{KL}(q_{\theta}(\mathbf{z}|\mathbf{x})||p_{\theta}(\mathbf{z}))+\mathbb{E}_{q_{\theta}(\mathbf{z}|\mathbf{x})}[\log p_{\theta}(\mathbf{x}|\mathbf{z})]$, where $\mathtt{KL}(q_{\theta}(\mathbf{z}|\mathbf{x})||p_{\theta}(\mathbf{z}))$ is the KL divergence between $q_{\theta}(\mathbf{z}|\mathbf{x})$ and $p_{\theta}(\mathbf{z})$, and $\mathbb{E}_{q_{\theta}(\mathbf{z}|\mathbf{x})}[\log p_{\theta}(\mathbf{x}|\mathbf{z})]$ is computed by the Stochastic Gradient Variational Bayes estimator [15]. The challenges for VAEs include undesirable stable equilibrium, Blurriness and so on [22].
II-C Autoregressive Models (AR Models)
Autoregressive (AR) models factorize $p_{\theta}(\mathbf{x})$ as a product of conditional probabilities,
$p_{\theta}(\mathbf{x})=p(x_{1},\ldots,x_{d})=\prod_{i=1}^{d}p_{\theta}(x_{i}\mid x_{1},\ldots,x_{i-1}),$ (3)
where $d$ is the sequence length. Such factorisation simplifies multivariate density estimation and has been widely adopted to model pixels sequentially in images [23, 24, 25].
To reduce the quadratic cost of attention in standard transformer-based AR models, several non-transformer architectures have recently been introduced. RWKV [26], Mamba [27], and RetNet [28] replace or augment attention with recurrent or state-space mechanisms. RWKV and Mamba employ purely recurrent designs that maintain a fixed-size memory, delivering linear-time inference over moderate sequence lengths but still facing challenges at extreme context sizes. RetNet updates hidden states through a retention mechanism, providing an efficient alternative to global self-attention. Although these architectures show promising results on language and other sequence tasks, their use as backbones for deep generative models is still limited. Future work that integrates them into generative pipelines may improve the trade-off among sample quality, scalability, and memory usage.
II-D Normalizing Flows (NFs)
@@
where $g^{-1}(\cdot)$ is the inverse of $g(\cdot)$, and $\mathtt{J}(g(\mathbf{z}))=\frac{\partial g}{\partial\mathbf{z}}$ is the Jacobian of $\mathbf{g}$. NFs construct more complicated non-linear invertible functions by compositing a set of $N$ bijective functions and define $g=g_{N}\circ\dots\circ g_{1}$.
II-E Diffusion Models
@@
where $Z_{\theta}=\int\exp(-s_{\theta}(\mathbf{x}))d\mathbf{x}$ is the normalization term. Since $Z_{\theta}$ is intractable to compute, the diffusion models learn the score function $\nabla_{\mathbf{x}}\log p_{\theta}(\mathbf{x})=-\nabla_{\mathbf{x}}s_{\theta}(\mathbf{x})$ instead.
@@
III Paradigms
This section presents methods for simulating the real world from the perspective of data dimension growth. It begins with 2D generation (Section III-A) for appearance modeling and then move to video generation (Section III-B) and 3D generation (Section III-C) by incorporating dynamics and geometry dimensions. Finally, by integrating all these three dimensions, recent advances in 4D generation (Section III-D) are presented.
III-A 2D Generation
@@ III-A1 Algorithms
DeepFloyd IF [37] is celebrated for its exceptional photorealism and advanced language understanding. This system is modular, featuring a static text encoder and three sequential pixel diffusion modules. Initially, the base model creates $\times64$ pixel images from textual descriptions. These images are then enhanced to $\times256$ pixels and further to $\times1024$ pixels by two super-resolution models. Each phase utilizes a static text encoder derived from the T5 [32] transformer to generate text embeddings, which are subsequently processed by a U-Net architecture with integrated cross-attention and attention pooling mechanisms.
Stable Diffusion (SD) [38], also known as Latent Diffusion Model (LDM), enhances training and inference efficiency on limited computational resources while producing high-quality and diverse images. The denoising process takes place in the latent space of pre-trained autoencoders, which map images into a spatial latent space. The underlying U-Net architecture is augmented with a cross-attention mechanism to model the conditional distribution, which can include text prompts, segmentation masks, and more. It used CLIP [30] text embeddings as condition and trained on the LAION [39] dataset to generate images at a resolution of $\times512$ (with a latent resolution of $\times64$). Building on Stable Diffusion, SDXL [40] employs a U-Net backbone that is three times larger. It introduces additional attention blocks and a larger cross-attention context by utilizing a second text encoder. Additionally, SDXL includes a refinement model that enhances the visual fidelity of samples generated by SDXL through a post-hoc image-to-image technique.
@@
III-B Video Generation
Figure 2: An illustration of video generation paradigm. Video generation models are constructed on top of image generation models by adding temporal layers or from scratch.
Due to the structural similarities between images and videos, early approaches to video generation predominantly adapted and fine-tuned existing 2D image generation models (Sec. III-A). The initial perspective on this challenge focused on incorporating mechanisms for temporal dynamic modeling, often by adding temporal layers (Fig. 2), such as attention and convolution, to the architecture. A common training strategy also involved mixing 2D images and video data to enhance the visual quality of the generated videos. Inspired by Sora [13], state-of-the-art models now frequently employ a diffusion transformer architecture. These models operate on a compressed spatio-temporal latent space by breaking down video into a series of “patches” that serve as tokens for the transformer. This approach processes both spatial and temporal information simultaneously. Despite their sophistication, these models typically retain the ability to process single images as one-frame videos, allowing them to leverage the vast data available for 2D image generation.
In this section, we classify these models into three main categories based on their underlying generative machine learning architectures. Figure 3 summarizes recent text-to-video generation techniques. For readers seeking a more in-depth exploration, detailed surveys are available [42, 43].
III-B1 Algorithms
@@
Evaluation. The evaluation of video generation models has evolved with the task’s increasing complexity. Early approaches relied on distribution-based metrics, most notably the Fréchet Video Distance (FVD) [83, 56, 50]. As a temporal extension of the Fréchet Inception Distance (FID) [84] for images, FVD compares spatiotemporal feature distributions to assess visual quality and coherence. More recent benchmarks, such as VBench [85], offer more granular analysis of specific attributes like motion smoothness and subject identity using features from models like CLIP [30] and DINO [86]. However, because automated metrics often misalign with human perception, the field has increasingly shifted towards human studies for a more holistic and accurate assessment, particularly for advanced open-domain models. Tab. II presents a human preference evaluation of modern video generation models.
Dimension ($\uparrow$) Hunyuan [87] Mochi [88] SoRA [13] Wan-14B [89]
Motion Quality 0.413 0.420 0.482 0.415
Human Artifacts 0.734 0.622 0.786 0.691
Pixel Stability 0.983 0.981 0.952 0.972
ID Consistency 0.935 0.930 0.925 0.946
Physical 0.898 0.728 0.933 0.939
Smoothness 0.890 0.530 0.930 0.910
Image Quality 0.605 0.530 0.665 0.640
Scene Quality 0.373 0.368 0.388 0.386
Stylization 0.386 0.403 0.606 0.328
Object Accuracy 0.912 0.949 0.932 0.952
TABLE II: Quantitative comparison between different text-to-video methods. Results are obtained from [89].
III-B2 Applications
@@
(3) Human animation in videos has gained significance in video generation, which plays a pivotal role in the world simulator, as discussed in Sec. III-B1. This is particularly important because humans are the most essential participants in the real world, making their realistic simulation crucial. Thanks to the early success in generative models, there are some representative works [46, 103, 104] that introduce generative adversarial networks (GAN) [14] for animating humans in videos. Despite these progress, the most essential problem in human video animation is the visual fidelity of the generated videos. ControlNet [105] and HumanSD [106] are plug-and-play methods for animating humans referring to poses based on a foundation text-to-image model, like Stable Diffusion [38]. Additionally, to resolve the generalization issue of these methods, animate-anyone [107] proposes a ReferenceNet to maintain more spatial details of the reference video and pushes generation quality in the wild to a new milestone. There are some follow-up works [108, 109] of animate-anyone trying to simplify the training architecture and costs. Besides, as the in-depth study of geometry and textures in the computer graphics area, there are some works that introduce 3D modeling into human video animation. Liquid Warping GAN [110], CustomHuman [111] and LatentMan [112] are early attempts introducing 3D human prior into the generation loop. The latest progress, MIMO [113], explicitly models the charter, 3D motion, and scene respectively to drive human animation in the wild. These methods with or without 3D priors make a great step to introduce humans into the loop of the world simulator.
@@
III-C 3D Generation
3D generation focuses on both geometry and appearance to better simulate real-world scenarios. In this section, we explore various 3D representations and generation algorithms, providing a structured overview of recent advancements. Specifically, we categorize 3D generation methods based on their input modalities, including Text-to-3D Generation, which directly synthesizes 3D content from textual descriptions, Image-to-3D Generation, which introduces image constraints to refine text-driven outputs, and Video-to-3D Generation, which leverages video priors for more consistent 3D generation. A chronological summary of these advancements is presented in Figure 7, while Table IV offers a comprehensive comparison of cutting-edge methods. Notably, several approaches span multiple categories, demonstrating the versatility of modern 3D generation techniques.
Instead of constructing 3D generation models from scratch, most existing approaches are highly coupled with 2D and video generation models to leverage their powerful appearance modeling capability for better 3D generation, as illustrated in Figs. 5, 8, and 10. First, the image prior encoded in 2D and geometry cues encoded in video generation models can be employed to provide supervision for 3D generation models. Second, 2D and video generation models can be fine-tuned to take additional 3D information (e.g., normals) as input to synthesize 3D-aware multi-view images to facilitate 3D generation.
III-C1 3D Representation
@@
III-C2 Algorithms
Figure 5: Comparison between different text-to-3D generation paradigms. Image generation models serve as critical components to construct text-to-3D approaches.
(1) Text-to-3D Generation. To generate 3D content from the text prompt by simulating the geometrics in the real world, extensive studies have been conducted and can be divided into three branches. Readers can refer to [128, 129, 130] for a more comprehensive survey in this field. Comparison of different branches of methods is illustrated in Fig. 5. As we can see, image generation models serve as critical components in text-to-3D approaches to provide supervision (i.e., SDS loss) or synthesize multi-view images for more accurate 3D generation.
[…]
Time ($\downarrow$) Quality ($\uparrow$) Alignment ($\uparrow$)
Magic3D [144] 40min 38.7 35.3
Fantasia3D [146] 45min 29.2 23.5
DreamFusion [142] 30min 24.9 24.0
ProlificDreamer [147] 240min 51.1 47.8
MVDream [143] 30min 53.2 42.3
DreamGaussian [150] 7min 19.9 19.8
VP3D [149] - 54.8 52.2
GaussianDreamer [157] - 54.0 -
TABLE III: Quantitative comparison between different text-to-3D methods. Results are obtained from [157].
Figure 6: Qualitative comparison between different text-to-3D methods.
Evaluation. Quantitative evaluation of text-to-3D methods remains an open problem. For subjective quality evaluation, common reference-based metrics (e.g., PSNR) are not applicable as groundtruth data are unavailable while non-reference quality (e.g., FID) metrics may not always align with human preferences. Consequently, most methods employ CLIP Score and CLIP R-Precision to evaluate the alignment of 3D models with text prompts. Recently, several benchmarks [158, 159] are established to comprehensively assess text-to-3D generation methods. Here, we report the quantitative scores of representative methods in Table III and present their visual results in Fig. 6. Readers can refer to [158, 159] for more details.